# Densities Of Primes And Primitive Roots

### v10

@@ Abstract
Let $u\neq\pm 1,v^{2}$ be a fixed integer, let $p\geq 2$ be a prime, and let $\operatorname{ord}_{p}(u)=d\>|\>p-1$ be the order of $u\operatorname{mod}p$. This note provides an effective lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-1}$ for the number of primes $p\leq x$ with a fixed primitive root $u\text{ mod }p$ for all large numbers $x\geq 1$. The current results in the literature have the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$, and restrictions on the fixed primitive root to a subset of at least three or more integers. August 30, 2025
@@ 1 Introduction
see [21]. In both of these results, the lower bound for the number of primes $p\leq x$ with a fixed primitive root in either of the subset $\mathcal{A}(q,r,s)$ or $\mathcal{B}(q,r,s)$ has the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$ for all large number $x\geq 1$.
@@ 3.3 Divisors Free Characteristic Function
(i) As the index $n\geq 1$ ranges over the integers relatively prime to p1p-1, the element $\tau^{n}\in\mathbb{F}_{p}$ ranges over the primitive roots $\text{mod }p$. Ergo, the equation
$\tau^{n}-u=0$ (3.4)
has a solution if and only if the fixed element $u\in\mathbb{F}_{p}$ is a primitive root. Next, replace $\psi(z)=e^{i2\pi kz/p}$ to obtain
@@
This follows from the geometric series identity $\sum_{0\leq k\leq N-1}w^{k}=(w^{N}-1)/(w-1)$ with $w\neq 1$, applied to the inner sum. For (ii), use the equation $\tau^{n}\bar{u}=1$, where $\bar{u}$ is the inverse of $u$, and apply the geometric series identity. ∎
4 Estimates Of Exponential Sums
This section provides simple estimates for the exponential sums of interest in this analysis. There are two objectives: To determine an upper bound, proved in Theorem 4.2, and to show that the finite Fourier transform
$\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{b}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=-\sum_{\gcd(n,p-1)=1}e^{i2\pi\tau^{n}/p}+E(p),$ (4.1)
where $E(p)$ is an error term, this is proved in Lemma 4.1. These are indirectly implied by the equidistribution of the subsets
$\{\tau^{n}:\gcd(n,p-1)=1\}=\{b\tau^{n}:\gcd(n,p-1)=1\}\subset\mathbb{F}_{p},$ (4.2)
for any $0\neq b\in\mathbb{F}_{p}$. The proofs of these results are entirely based on established results and elementary techniques.
4.1 Incomplete And Complete Exponential Sums
Let $f:\mathbb{C}\longrightarrow\mathbb{C}$ be a function, and let $q\in\mathbb{N}$ be a large integer. The finite Fourier transform
$\hat{f}(t)=\frac{1}{q}\sum_{0\leq s\leq q-1}e^{i\pi st/q}$ (4.3)
and its inverse are used here to derive a summation kernel function, which is almost identical to the Dirichlet kernel.
Definition 4.1.
Let $p$ be a prime and let $\omega=e^{i2\pi/p}$ be a root of unity. The finite summation kernel is defined by the finite Fourier transform identity
This simple identity is very effective in computing upper bounds of some exponential sums
$\sum_{n\leq x}f(n)=\sum_{n\leq x}\mathcal{K}(f(n)),$ (4.5)
where $x\leq p$. Two applications are illustrated here.
Theorem 4.1.
([43], [31]) Let $p\geq 2$ be a large prime, and let $\tau\in\mathbb{F}_{p}$ be an element of large multiplicative order $\operatorname{ord}_{p}(\tau)\mid p-1$. Then, for any $b\in[1,p-1]$, and $x\leq p-1$,
$\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}\ll p^{1/2}\log p.$ (4.6)
@@ 4.1 Incomplete And Complete Exponential Sums
Let $p$ be a large prime, and let $f(n)=e^{i2\pi b\tau^{n}/p}$, where $\tau$ is a primitive root modulo $p$. Applying the finite summation kernel in Definition 4.1, yields
$\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}=\sum_{n\leq x}\frac{1}{p}\sum_{0\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{i2\pi b\tau^{s}/p}.$ (4.7)
Observe that the term $t=0$ contributes x/p-x/p, and rearranging it yield
$\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}$ $=$ $\frac{1}{p}\sum_{n\leq x,}\sum_{1\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{i2\pi b\tau^{s}/p}-\frac{x}{p}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq q-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{i2\pi b\tau^{s}/p}\right)\left(\sum_{n\leq x}\omega^{tn}\right)-\frac{x}{p}.$
Taking absolute value, and applying Lemma 4.2, and Lemma 4.4, yield
$\left|\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}\right|$ $\leq$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left|\sum_{0\leq s\leq p-1}\omega^{-ts}e^{i2\pi b\tau^{s}/p}\right|\cdot\left|\sum_{n\leq x}\omega^{tn}\right|+\frac{x}{p}$
$\ll$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(p^{1/2}\log p\right)\cdot\left(\frac{2p}{\pi t}\right)+\frac{x}{q}$
$\ll$ $p^{1/2}\log^{2}p.$
The last summation in (4.1) uses the estimate
$\sum_{1\leq t\leq p-1}\frac{1}{t}\ll\log p$ (4.10)
and $x/p\leq 1$. ∎
This appears to be the best possible upper bound. The above proof generalizes the sum of resolvents method used in [31]. Here, it is reformulated as a finite Fourier transform method, which is applicable to a wide range of functions. A similar upper bound for composite moduli $p=m$ is also proved, [op. cit., equation (2.29)].
@@
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi b\tau^{n}/p}\ll p^{1-\varepsilon}$ (4.11)
for any $b\in[1,p-1]$, and any arbitrary small number $\varepsilon\in(0,1/2)$.
@@
see Definition 4.1. Use the inclusion exclusion principle to rewrite the exponential sum as
@@
$=$ $\sum_{n\leq p-1}\frac{1}{p}\sum_{1\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}}\sum_{\begin{subarray}{c}d\mid p-1\\ d\mid n\end{subarray}}\mu(d)-\frac{\varphi(p)}{p}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi b\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p)}{p}.$
@@
$\leq$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left|\sum_{1\leq s\leq p-1}\omega^{-ts}e^{i2\pi b\tau^{s}/p}\right|\cdot\left|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right|+\frac{\varphi(p)}{p}$
$\ll$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(2p^{1/2}\log p\right)\cdot\left|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right|+\frac{\varphi(p)}{p}$
$\ll$ $\frac{1}{p}\left(2p^{1/2}\log p\right)\cdot\sum_{1\leq t\leq p-1}\left|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right|+\frac{\varphi(p)}{p}$
@@
where $\delta>0$ is a small number. This is restated in the simpler notation $p^{1/2+\delta}(\log p)^{2}\leq p^{1-\varepsilon}$ for any arbitrary small number $\varepsilon\in(0,1/2)$. ∎
The upper bound given in Theorem 4.2 seems to be optimum. A different proof, which has a weaker upper bound, appears in [14, Theorem 6], and related results are given in [5], [13], [19], [3], [2], [4], [14], [12] and [20, Theorem 1].
4.2 Finite Fourier Transforms and Equivalent Exponential Sums
An asymptotic relation for the finite Fourier transform of the exponential sum
$V(1,b)=\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi b\tau^{n}/p},$ (4.16)
in terms of simpler exponential sum up to an error term, is provided here.
Lemma 4.1.
Let $p\geq 2$ be a large prime. If $\tau$ be a primitive root modulo $p$, then
$\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{b}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}+O(p^{1/2+\delta}(\log p)^{2}),$ (4.17)
@@ 4.2 Finite Fourier Transforms and Equivalent Exponential Sums
For $b\geq 1$, the exponential sum has the representation
$V(1,b)$ $=$ $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}$
@@
$V(1,1)$ $=$ $\sum_{\gcd(n,p-1)=1}e^{\frac{i2\pi\tau^{n}}{p}}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p-1)}{p},$
@@
$V(1,b)-V(1,1)$ $=$ $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}$
@@
Let $a\in[1,p-1]$. Taking the finite Fourier transform of the difference D(a,b)=V(1,b)V(1,1)D(a,b)=V(1,b)-V(1,1) returns
$\widehat{D(a,b)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\left(\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}\right)$
$=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\times\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi b\tau^{s}}{p}}-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\right)$
$\times\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}\sum_{1\leq b\leq p-1}e^{\frac{i2\pi b(\tau^{s}-a)}{p}}-\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\right)$
Now at $a=1$, $\tau^{s}-a\neq 0$ for any $s\in[1,p-1]$. Thus, using the geometric sum identity $\sum_{1\leq u\leq p-1}e^{i2\pi u/p}=-1$ to simplify the last expression yields
$\widehat{D(1,b)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{b}{p}}\left(\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}\right)$
@@
$\widehat{V(1,b)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{b}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}$
@@
$\sum_{1\leq t\leq p-1}\Bigg{|}\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\Bigg{|}$ $=$ $\sum_{1\leq t\leq p-1}\Bigg{|}\sum_{\gcd(n,p-1)=1}\omega^{tn}\Bigg{|}$ (4.24)
@@
$\Bigg{|}1-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\Bigg{|}=\Bigg{|}1-\sum_{1\leq s\leq p-1}\chi(s)\psi(s)\Bigg{|}$ $\leq$ $2p^{1/2}\log p,$ (4.25)
where $\chi(s)=e^{i\pi st/p}$, and $\psi(s)=e^{i2\pi\tau^{s}/p}$. Taking absolute value of the remainder term in (4.2) and replacing (4.24), and (4.25), return
@@
The same proof works for many other subsets of elements $\mathcal{A}\subset\mathbb{F}_{p}$. For example,
$\sum_{n\in\mathcal{A}}e^{i2\pi b\tau^{n}/p}=\sum_{n\in\mathcal{A}}e^{i2\pi\tau^{n}/p}+O(p^{1/2+\delta}\log^{c}p),$ (4.27)
for some constant $c>0$.
4.3 Finite Summation Kernels And Gaussian Sums
@@
Let $p\geq 2$ be large prime and let $\omega=e^{i2\pi/p}$ be a $p$th root of unity, and let $t\in[1,p-1]$. Then,
  1. (i)
    $\bigg{|}\sum_{n\leq p-1}\omega^{tn}\bigg{|}\leq\frac{2p}{\pi t}.$
  2. (ii)
    $\sum_{1\leq t\leq p-1}\bigg{|}\sum_{n\leq p-1}\omega^{tn}\bigg{|}\leq 2p\log p.$

### v12

@@ Abstract
Let $u\neq\pm 1,v^{2}$ be a fixed integer, let $p\geq 2$ be a prime, and let $\operatorname{ord}_{p}(u)=d\>|\>p-1$ be the order of $u\operatorname{mod}p$. This note provides an effective lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-1}$ for the number of primes $p\leq x$ with a fixed primitive root $u\text{ mod }p$ for all large numbers $x\geq 1$. The current results in the literature have the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$, and restrictions on the fixed primitive root to a subset of at least three or more integers. October 8, 2025
@@ 1 Introduction
see [HB86]. In both of these results, the lower bound for the number of primes $p\leq x$ with a fixed primitive root in either of the subset $\mathcal{A}(q,r,s)$ or $\mathcal{B}(q,r,s)$ has the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$ for all large number $x\geq 1$.
@@ 3.3 Divisors Free Characteristic Function
(i) Set the additive character $\psi(s)=e^{i2\pi as/p}\in\mathbb{C}$. Now, as the index $n\in\mathscr{R}=\{n<p:\gcd(n,p-1)=1\}$ ranges over the integers relatively prime to $\varphi(p)=p-1$, the element $\tau^{n}\in\mathbb{F}_{p}^{\times}$ ranges over the primitive roots modulo $p$. Accordingly, the equation
$a=\tau^{n}-u=0$ (3.4)
has a unique solution $n\in\mathscr{R}$ if and only if the fixed element $u\in\mathbb{F}_{p}$ is a primitive root. This implies that the inner sum in
@@
collapses to $\sum_{0\leq s<p}e^{i2\pi as/p}=\sum_{0\leq s<p}1=p$. Otherwise, if the element $u\in\mathbb{F}_{p}$ is not a primitive root, then the equation (3.4) has no solution $n\in\mathscr{R}$, and the inner sum in (3.5) collapses to $\sum_{0\leq s<p}e^{i2\pi as/p}=0$, this follows from the geometric series formula $\sum_{0\leq n\leq N-1}w^{n}=(w^{N}-1)/(w-1)$, where $w=e^{i2\pi a/p}\neq 1$ and $N=p$. This completes the verification. ∎
4 Finite Summation Kernel and Gauss Sum
An upper bound for some elementary exponential sums are provided in this section.
4.1 Finite Summation Kernel
Lemma 4.1.
$\sum_{1\leq t\leq p-1}\Bigg|\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\omega^{tn}\Bigg|\ll p^{1+\delta}\log p,$ (4.1)
where $\delta>0$ is a small number.
@@ 4.1 Finite Summation Kernel
Use the inclusion exclusion principle to rewrite the exponential sum as
$=$ $\sum_{d\mid p-1}\mu(d)\sum_{m\leq(p-1)/d}\omega^{dtm}$
Now, the parameters are $p$ prime, $\omega=e^{i2\pi/p}$, the integers $t\in[1,p-1]$ and $d\leq p-1<p$. This data implies that $\pi dt/p\neq k\pi$ with $k\in\mathbb{Z}$, so the sine function $\sin(\pi dt/p)\neq 0$ is well defined. Consequently, the absolute value satisfies
$\left|\frac{\omega^{dt}-\omega^{dt((p-1)/d+1)}}{1-\omega^{dt}}\right|\leq\left|\frac{2}{\sin(\pi dt/p)}\right|.$ (4.3)
For each $d\mid p-1$, the map $t\longrightarrow z\equiv dt\bmod p$ is a permutation in the finite field $\mathbb{F}_{p}$. Thus, using standard manipulations, and $z/2\leq\sin(z)<z$ for $0<|z|<\pi/2$, the last expression becomes
$\leq$ $\sum_{d\mid p-1,}\sum_{1\leq z\leq p-1}\frac{2p}{\pi z}$
$\ll$ $p^{1+\delta}\log p,$
4.2 Gauss Sum
Some elementary exponential sums estimates are provided in this section.
Lemma 4.2.
(Gauss sums)Let $p\geq 2$ be a prime, let $\chi(t)=e^{i2\pi t/p}$ and $\psi(t)=e^{i2\pi\tau^{t}/p}$ be a pair of characters. Then, the Gaussian sum has the upper bound
$\left|\sum_{1\leq t\leq p-1}\chi(t)\psi(t)\right|\leq 2p^{1/2}\log p.$
5 Estimates of Power Exponential Sums
The estimate for the power sum with relatively prime index is based on the identity
$\frac{1}{p}\sum_{0\leq t\leq p-1,}\sum_{0\leq s\leq p-1}\omega^{t(n-s)}f(s)=f(n),$ (5.1)
where $\omega=e^{i2\pi/p}$ and $x\leq p-1$.
5.1 Power Exponential Sum with Relatively Prime Index
@@
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi b\tau^{n}/p}\ll p^{1/2+\delta}(\log p)^{2},$
where $\delta>0$ is a small real number and the implied constant is independent of $b\neq 0$.
@@
see (5.1). Use the inclusion exclusion principle to rewrite the exponential sum as
@@
$=$ $\sum_{n\leq p-1}\frac{1}{p}\sum_{1\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}}\sum_{\begin{subarray}{c}d\mid p-1\\ d\mid n\end{subarray}}\mu(d)-\frac{\varphi(p-1)}{p}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi b\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p-1)}{p}.$
@@
$\leq$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left|\sum_{1\leq s\leq p-1}\omega^{-ts}e^{i2\pi b\tau^{s}/p}\right|\cdot\left|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right|+\frac{\varphi(p-1)}{p}$
$\ll$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(2p^{1/2}\log p\right)\cdot\left|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right|+\frac{\varphi(p-1)}{p}$
$\ll$ $\frac{1}{p}\left(2p^{1/2}\log p\right)\cdot\sum_{1\leq t\leq p-1}\left|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right|+\frac{\varphi(p-1)}{p}$
@@
where $\delta>0$ is a small number. ∎
A different approach to this result appears in [14, Theorem 6], and related results are given in [13], [19], [5], and [20, Theorem 1]. The upper bound given in Theorem 5.1 seems to be optimum. A different proof, which has a weaker upper bound, appears in [14, Theorem 6], and related results are given in [5], [13], [20], and [20, Theorem 1].
5.2 FFT of Power Exponential Sum with Relatively Prime Index
For any fixed $0\neq b\in\mathbb{F}_{p}$, the map $\tau^{n}\longrightarrow b\tau^{n}$ is one-to-one (permutation) in $\mathbb{F}_{p}$. Consequently, the subsets
$\{\tau^{n}:\gcd(n,p-1)=1\}\quad\text{ and }\quad\{b\tau^{n}:\gcd(n,p-1)=1\}\subset\mathbb{F}_{p}$ (5.6)
have the same cardinalities. As a direct consequence the exponential sums
$V(b)=\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi b\tau^{n}/p}\quad\text{ and }\quad V(1)=\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\tau^{n}/p},$ (5.7)
have the same upper bound up to an error term. An asymptotic relation for the finite Fourier transform (FFT) of the exponential sums (5.7) is provided here.
Theorem 5.2.
Let $p\geq 2$ be a large prime. If $\tau$ be a primitive root modulo $p$ and $a<x=o(p)$ is not a primitive root, then
$\widehat{V(a)}=\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}+O(p^{1/2+\delta}(\log p)^{2}),$
@@ 5.2 FFT of Power Exponential Sum with Relatively Prime Index
For $a\in[1,x]$ and $b\in[1,p-1]$, the exponential sum has the representation
$V(b)$ $=$ $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}$
@@
$V(1)$ $=$ $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi a\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p-1)}{p},$
@@
$V(b)-V(1)$ $=$ $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}$
@@
Taking the finite Fourier transform of the difference D(b)=V(b)V(1)D(b)=V(b)-V(1) returns
$\widehat{D(a)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\left(\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}\right)$
$=$ $\frac{1}{p}\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{b}{p}}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi b\tau^{s}}{p}}-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi a\tau^{s}}{p}}\right)$
$\hskip 126.47249pt\times\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}\sum_{1\leq b\leq p-1}e^{\frac{i2\pi b(\tau^{s}-a)}{p}}\right$.
$\qquad-\left.\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\right)\times\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right).$
Now in the range $a<x=o(p)$, $\tau^{s}-a\neq 0$ for any $s\in\{s<p-1:\gcd(s,p-1)=1\}$. Thus, using the geometric sum identity $\sum_{1\leq u\leq p-1}e^{i2\pi au/p}=-1$ to simplify the last expression yields
$\widehat{D(a)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\left(\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}\right)$ (5.12)
@@
$\widehat{V(a)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}$
@@
$\sum_{1\leq t\leq p-1}\Bigg|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\Bigg|$ $=$ $\sum_{1\leq t\leq p-1}\Bigg|\sum_{\gcd(n,p-1)=1}\omega^{tn}\Bigg|$
@@
$\Bigg|1-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi a\tau^{s}}{p}}\Bigg|=\Bigg|1-\sum_{1\leq s\leq p-1}\chi(s)\psi(s)\Bigg|$ $\leq$ $2p^{1/2}\log p,$ (5.15)
where $\chi(s)=e^{i\pi st/p}$, and $\psi(s)=e^{i2\pi a\tau^{s}/p}$. Taking absolute value of the remainder term in (5.2) and replacing (5.2), and (5.15), return
@@
Let $p\geq 2$ be a large prime. If $\tau$ be a primitive root modulo $p$ and $0\neq a=o(p)$ is not a primitive root, then
$\Bigg|\widehat{V(a)}\Bigg|=\Bigg|\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}\Bigg|=O(p^{1/2+\delta}(\log p)^{2}),$
where $\delta>0$ is a small number and the implied constant is independent of $a$ and $b\in[1,p-1]$.